GlueX TOF Meeting, November 18, 2014

From GlueXWiki
Jump to: navigation, search

GlueX Time-of-Flight Meeting
Tuesday, November 18, 2014
11:00 am EST
JLab: CEBAF Center F326/327


  1. Announcements
  2. Work in the Hall (Aristeidis)
  3. Commissioning Plan Review
  4. Calibration Progress (Sasha)
  5. Equipment Loan



Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2014-4Q on the JLab CUE. This directory is accessible from the web at .



  • FSU: Hussein Al Ghoul, Paul Eugenio, Sasha Ostrovidov, Aristeidis Tsaris
  • JLab: Mark Ito (chair)

Commissioning Plan Review

Paul was prepared to present the commissioning run plan for the TOF at this morning's Run Coordination Meeting, but the agenda did not include this discussion. We went over the components of our request:

  1. For calibration we need straight tracks (no field) going through the TOF.
  2. We need a TOF-based trigger to illuminate the corners of the TOF array that are beyond the FCAL.
  3. We need running in two target positions:
    1. Nominal position to allow tracking with the FDC.
    2. Forward position to allow tracks to hit the corners of the TOF. In the nominal position parts of the TOF are shadowed by the solenoid iron.
  4. High-rate running to assess TOF performance under those conditions. For this we want as thick a target and as high a beam rate as possible.

FADC Voltage Range

We have been running with a full range of 1 Volt so far. All tests at FSU were done with the 2-V range. Sasha has reported 30% of the pulses with overflows in the last cosmic runs he took some weeks ago. Fernando Barbosa has questioned whether we need to make this change. It is possible that the cosmic data has much larger average pulse height than beam data due to the high number of vertical cosmics. Sasha will look at the beam data from last night to check whether we want to proceed with changing the FADC range.

TDC Timing

The upgrade to the firmware that we discussed at the last TOF meeting has still not been installed. Sascha Somov had told Paul that it would be in by now. In addition we have still not seen the trigger signal in the TDC data. We need to continue to pursue this issue.

Sasha O. is concerned that there might be timing jitter from board-to-board in the system due to the trigger pulse being presented to different boards at different times. Mark stated that this should not happen; if it occurs in the data then we need to address it, but we should push ahead on the assumption that there are no random time shifts between any two channels in the CAEN TDC crate.


We have transitioned to pursuing beam-based calibration only.

The last run examined, run 995, had no TDC data for the TOF. The read-out window was not adjusted when going from the stand-alone TOF trigger to the global FCAL trigger. Last Thursday Paul adjusted the window. Sasha O. will look at the data taken yesterday and see if the TDC data is present.

There is still confusion about the pulse-integral data. The behavior of the data seems different than that of the raw mode data. The effect of NSB/NSA on the integral data is not clear. Mark will try to find the document describing the algorithm used by the firmware. It would be helpful to have data with the full raw pulse and the pulse-integral summary in the same event so an offline check can be made.

Equipment Loan

Paul is looking through the items and will return those that are not needed for reading out the mini-TOF.

Action Items

  1. Look at overflow fraction in beam-based FADC data.
  2. Request the TS change to synch the CAEN TDC system clock to the global system clock.
  3. Look for TDC hits in the read-out window of beam data.
  4. Get the document on the FADC pulse-integral algorithm.
  5. Ask for raw and summary FADC data in the same event.
  6. Assess need for loaned equipment.